3. The basics of the Capital Asset Pricing Model Aa Aa which of the following are assumptions of the Capital Asset Pricing Model (CAPM)? Check all that apply Investors have identical estimates of expected retums but not of variances. Investors can borrow an unlimited amount at a risk-free rate Investors assume that their investment activities wont affect the price of a stock. All investors are price givers. Consider the equation for the Capital Asset Pricing Model (CAPM): Cov(, M) In this equation, the term rN R) represents the Suppose that the markets average excess retu on stocks is 6.00% and that the risk-free rate is 2.00%. Complete the following table by computing expected retums to stocks for each beta coefficient using the Capital Asset Pricing Model (CAPM): bi -0.30 0.50 1.00 5.00 Return to stocks (%) 8.0D 5.00 Based on the CAPM and your calculations for the retum to stocks, what does it mean when the coefficient b 17 O The stock is more volatile than the market. O The stock is less volatile than the arket. The stocks retum oorrelates with the stock market as a whole.
Answer #1

First part of Question is What is Rm - RF in CAPM Equation.

Rm is the Rate of return of Market and Rf is Risk Free rate of Return.

Rm - Rf therefore is excess return of market over risk free rate.

Second Part involves calculation of Return on Stocks with different Beta

Rf = 2%

Rm - Rf = 6%

When Beta = 1%

Formula to calculate return = Rf + Beta (Rm-Rf)

Return = 2 + 1 (6)

= 8%

When Beta = 5%

Return = 2 +5 (6)

= 2 + 30

32%